OpenStudy (anonymous):

In factored form n^2(n+8)-4(n+8) Medal+fan :)

OpenStudy (freckles):

factor out the (n+8)

OpenStudy (freckles):

recall distributive property a(b+c)=ab+ac this can also be used the other way ab+ac=a(b+c) image the a here being (n+8) and the b being n^2 and the c being -4